# LLMCheckup: Conversational Examination of Large Language Models via Interpretability Tools

### v1

@@ Abstract
Interpretability tools that offer explanations in the form of a dialogue have demonstrated their efficacy in enhancing users’ understanding Slack et al. (2023); Shen et al. (2023), as one-off explanations may occasionally fall short in providing sufficient information to the user. Current solutions for dialogue-based explanations, however, require many dependencies and are not easily transferable to tasks they were not designed for. With LLMCheckup [^1] Code: https://github.com/DFKI-NLP/LLMCheckup Demo: https://youtu.be/ZwN8ZQSXoOU, we present an easily accessible tool that allows users to chat with any state-of-the-art large language model (LLM) about its behavior. We enable LLMs to generate all explanations by themselves and take care of intent recognition without fine-tuning, by connecting them with a broad spectrum of Explainable AI (XAI) tools, e.g. feature attributions, embedding-based similarity, and prompting strategies for counterfactual and rationale generation. LLM (self-) explanations are presented as an interactive dialogue that supports follow-up questions and generates suggestions. LLMCheckup provides tutorials for operations available in the system, catering to individuals with varying levels of expertise in XAI and supports multiple input modalities. We introduce a new parsing strategy called multi-prompt parsing substantially enhancing the parsing accuracy of LLMs. Finally, we showcase the tasks of fact checking and commonsense question answering.
@@
via Interpretability Tools
Qianli Wang1,2   Tatiana $*{}^{*}$1,3   Nils $*{}^{*}$1 Josef van Genabith1,3   Leonhard Hennig1   Sebastian Möller1,2 ${}^{1}$erman Research Center for Artificial Intelligence (DFKI) ${}^{2}$echnische Universität Berlin, Germany ${}^{3}$aarland Informatics Campus, Saarbrücken, Germany {firstname.lastname}@dfki.de
@@ 1 Introduction
By contrast, our system, LLMCheckup, only requires a single LLM and puts it on “quadruple duty”: (1) Analyzing users’ (explanation) requests (§2.1, §5.2), (2) performing downstream tasks (§4), (3) providing explanations for its outputs (§3), and (4) responding to the users in natural language (§2.3). Instead of using many LMs to explain the behavior of another LLM, LLMCheckup allows us to directly request the same LLM used for parsing to self-explain its own behavior. Any LLM can be deployed “out of the box” and confronted with diverse NLP tasks, explanation generation and intent recognition. LLMCheckup provides a unified and compact framework, effectively consolidating and obviating the requirements for additional LMs to generate explanations.
@@ 2 LLMCheckup
LLMCheckup is an interface for chatting with any LLM about its behavior. We connect several interpretability tools (§3), s.t. self-explanations take model internals, datasets and documentation into account. User requests for explanations are recognized via a text-to-SQL task performed by the LLM (§2.1).
@@ 2.1 System architecture
Figure 2 illustrates the interaction flow of LLMCheckup. When a user asks a question on the user interface, the question will be parsed as an SQL-like query by the LLM in use. E.g., the first user question in Figure 1 would be parsed as "filter id 26 and rationalize". The corresponding operation (see Table 6) will then be matched (parsing) and executed. Finally, the generated explanation will be filled into a template that will be returned back to the user Slack et al. (2023); Feldhus et al. (2023) (response generation).
Parsing
To enhance the recognition of users’ intentions, we employ two methods: Guided Decoding and Multi-prompt Parsing.
@@
Guided Decoding (GD) ensures that the generated output adheres to predefined grammatical rules and constraints Shin et al. (2021) and that generated parses align with predefined operation sets Slack et al. (2023). GD is generally more suitable for smaller LMs.
@@
As an alternative to GD, we also propose and implement a novel Multi-prompt Parsing (MP) approach. While GD pre-selects prompts based on the embedding similarity with user input, the model does not see all the available operations at once and the pre-selection may not include the examples for the actual operation required. With MP, we test whether showing all possible operations in a simplified format as the first step (i.e., without any attributes such as the instance ID or the number of samples) and then additionally prompting the model to fill in more fine-grained attributes can improve performance. We provide further details about MP in Appendix B. §5.2 evaluates the performance of both parsing approaches.
@@ 2.2 Key features
We identify five auto-regressive LLMs representative of the current state-of-the-art in open-source NLP (as indicated in the left column of Table 2) available through Hugging Face Transformers Wolf et al. (2020). The diverse choice of models demonstrates that our system is generalizable and that in principle it is theoretically possible to choose any Transformer-type model to be explained. While Falcon-1B Penedo et al. (2023) and Pythia-2.8B Biderman et al. (2023) are available for users with limited hardware resources (RAM/GPU), it is generally not recommended to use them due to their small model size, which may negatively affect the user’s perception of the performance. Llama2-7B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) are both mid-sized with 7B parameters, while Stable Beluga 2 Mukherjee et al. (2023) is a fine-tuned version of Llama2-70B. To facilitate the deployment of large models in a local environment, LLMCheckup offers support for various forms of LLMs. This includes the usage of quantized models through GPTQ Frantar et al. (2023), loading models in 4-bits with the assistance of bitsandbytes Dettmers et al. (2022), and the implementation of a peer-to-peer solution using Petals Borzunov et al. (2023), enabling efficient deployment on a custom-level GPU.
@@
To help laypeople get background knowledge in XAI, we introduce a tutorial functionality. It is based on prompting with different roles corresponding to levels of expertise in XAI (Figure 3) and enables us to provide tailored meta-explanations of supported operations to individuals. For example, for the beginner level, we add a system prompt hinting at the expertise: “As a NLP beginner, could you explain me the question: What is data augmentation?” (Figure 4).
@@
LLMs may sometimes generate responses that appear plausible but are actually incorrect predictions known as hallucinations Welleck et al. (2020). LLMCheckup provides users with the capability to access information by conducting search through external knowledge bases, promoted by the integration of googlesearch [^2] https://github.com/Nv7-GitHub/googlesearch (Figure 5). Users can cross-reference the retrieved information with the provided explanations, thereby achieving a more comprehensive understanding.
Customized inputs & prompts
In comparison to TalkToModel Slack et al. (2023), which was limited to three datasets, LLMCheckup offers users the freedom to enter custom inputs according to their preferences, going beyond just querying instances from the provided dataset. In addition, inspired by PromptSource Bach et al. (2022), users can insert pre-defined prompts to the explanation step (§3.2), but the Prompt Editor also allows fully customized prompts, allowing for a personalized and tailored experience. All custom inputs will be saved and can be inspected and reused later on via a dedicated custom input history viewer.
@@
Motivated by Malandri et al. (2023), LLMCheckup not only accepts text input from users but also provides support for other modalities such as images and audio. To facilitate this, we employ specific packages and models tailored to each modality. For optical character recognition (OCR), we use EasyOCR [^3] https://github.com/JaidedAI/EasyOCR. For audio recognition, we employ a lightweight fairseq S2T [^4] https://huggingface.co/facebook/s2t-small-librispeech-asr model Wang et al. (2020) trained on Automatic Speech Recognition (ASR).
Suggestion of follow-up questions
To guide the user through the conversation, we implemented a suggestions mode. When activated, the user will get suggestions for related operations that LLMCheckup can perform (see Appendix C for more detail), e.g., if the user asks about the top $k$ attributed tokens for a specific sample, they will get a suggestion to have a look at the rationales generated by the model since both token attributions and rationalizations belong to the “Explanation" category also displayed in the user interface. Additionally, for each generated suggestion we check whether it already appears in the dialogue history to make sure that the user does not receive repetitive suggestions for the operations that they have already been performed. E.g., if the user inquires about the counterfactual operation and the model explains how it works, LLMCheckup will store this information and will not suggest explaining counterfactuals again.
@@
2.3 Interface
LLMCheckup provides a user interface including a chat window to enter questions and settings and is implemented in Flask and can be run as a Docker container. LLMCheckup builds on top of the chat window from TalkToModel Slack et al. (2023) for continuous interaction as well as buttons for question generation and the dataset viewer from InterroLang Feldhus et al. (2023), which facilitates dataset exploration and generates specific questions for all available operations to inspire users to come up with their own questions. Figure 3 shows the LLMCheckup interface.
In comparison to TalkToModel Slack et al. (2023) and InterroLang Feldhus et al. (2023), our interface provides the following new components:
  • Checkbox for selecting XAI user expertise level;
  • Input field for custom inputs;
  • Prompt editor including a checkbox for selecting additional prompts and a display of the currently used system prompt;
  • Support for multimodal input through buttons for image upload and audio recording.
Dataset COVID-Fact ECQA
Mistral 7B 0.66 0.88 0.96 0.50 0.76
Llama2 7B 0.65 0.88 0.94 0.50 0.76
Stable Beluga 2 70B 1.00 0.85 0.96 1.00 0.73
@@ 3 NLP explainability tools
Table 3 and Table 4 show examples of explanations for each supported explainability method by LLMCheckup.
@@ 3.1 White-box
We provide white-box explainability operations that are gradient-based or embedding-related, both of which are not easily applicable to the prompting paradigm of LLMs.
@@
Feature attribution methods quantify the contribution of each input token towards the final outcome. In LLMCheckup, we deploy various auto-regressive models (§2.2), for which Inseq Sarti et al. (2023) is used to determine attribution scores. We support representative methods from Inseq, including Input x Gradient Simonyan et al. (2014), Attention Bahdanau et al. (2015), LIME Ribeiro et al. (2016), and Integrated Gradients Sundararajan et al. (2017) [^5] Details on the Inseq integration are described in App. E..
@@ 3.2 Black-box
Augmentation involves synthesizing new instances by replacing text spans of the input while preserving the semantic meaning and predicted outcomes Ross et al. (2022). As a baseline, NLPAug is used to substitute input by WordNet’s Miller (1995) synonyms [^6] https://github.com/makcedward/nlpaug NLPAug offers various techniques for generating augmented text, including back-translation, word embedding substitution, etc.. However, we constrain the augmentation to only utilize WordNet to avoid additional LMs.. Alternatively, data augmentation can be achieved by providing the LLM with a few demonstrations as prompts Dai et al. (2023).
@@
Unlike data augmentation, counterfactuals are about input edits causing the predicted outcome to be different Wu et al. (2021); Chen et al. (2023). Counterfactuals are generated by prompting LLMs with manually crafted demonstrations.
@@ 4 Use cases
In this paper, we demonstrate the usefulness of LLMCheckup with two typical NLP tasks: Fact checking and commonsense question answering.
@@ 4.1 Fact checking
The importance of fact checking has grown significantly due to the rapid dissemination of both accurate information and misinformation within the modern media ecosystem Guo et al. (2022). COVID-Fact Saakyan et al. (2021) is a dataset resembling FEVER Thorne et al. (2018) that encompasses various claims, supporting evidence for those claims, and contradictory claims that have been debunked by the presented evidence.
@@ 4.2 Commonsense question answering
Unlike question answering, commonsense question answering (CQA) involves the utilization of background knowledge that may not be explicitly provided in the given context Ostermann et al. (2018). The challenge lies in effectively integrating a system’s comprehension of commonsense knowledge and leveraging it to provide accurate responses to questions. ECQA Aggarwal et al. (2021) is a dataset designed for CQA. Each instance in ECQA consists of a question, multiple answer choices, and a range of explanations. Positive explanations aim to provide support for the correct choice, while negative explanations serve to refute incorrect choices. Additionally, free-text explanations in natural language are included as general justifications.

### v2

@@ Abstract
Interpretability tools that offer explanations in the form of a dialogue have demonstrated their efficacy in enhancing users’ understanding Slack et al. (2023); Shen et al. (2023), as one-off explanations may fall short in providing sufficient information to the user. Current solutions for dialogue-based explanations, however, often require external tools and modules and are not easily transferable to tasks they were not designed for. With LLMCheckup111 https://github.com/DFKI-NLP/LLMCheckup, we present an easily accessible tool that allows users to chat with any state-of-the-art large language model (LLM) about its behavior. We enable LLMs to generate explanations and perform user intent recognition without fine-tuning, by connecting them with a broad spectrum of Explainable AI (XAI) methods, including white-box explainability tools such as feature attributions, and self-explanations (e.g., for rationale generation). LLM-based (self-)explanations are presented as an interactive dialogue that supports follow-up questions and generates suggestions. LLMCheckup provides tutorials for operations available in the system, catering to individuals with varying levels of expertise in XAI and supporting multiple input modalities. We introduce a new parsing strategy that substantially enhances the user intent recognition accuracy of the LLM. Finally, we showcase LLMCheckup for the tasks of fact checking and commonsense question answering.
@@
via Interpretability Tools and Self-Explanations
Qianli Wang1,2   Tatiana Anikina1,3   Nils Feldhus1 Josef van Genabith1,3   Leonhard Hennig1   Sebastian Möller1,2 1German Research Center for Artificial Intelligence (DFKI) 2Technische Universität Berlin, Germany 3Saarland Informatics Campus, Saarbrücken, Germany {firstname.lastname}@dfki.de
@@ 1 Introduction
By contrast, our framework, LLMCheckup, only requires a single LLM and puts it on “quadruple duty”: (1) Analyzing users’ (explanation) requests (§2.1, §5.1), (2) performing downstream tasks (§4), (3) providing explanations for its outputs (§3), and (4) responding to the users in natural language (§2.3). Instead of using many different LMs to explain the behavior of another LLM, LLMCheckup allows us to directly employ the same LLM used for user intent recognition to self-explain its own behavior. The advantage of a single-model approach is that it simplifies the engineering aspect of building an XAI system without multiple external modules and separately fine-tuned models. At the same time, we consistently achieve good performance even with a single model, as modern LLMs are very powerful and can handle a wide range of tasks including user intent recognition and explanation generation. Thus, LLMCheckup provides a unified and compact framework that is useful for future user studies in the context of human-computer interaction and explainability.
@@ 2 LLMCheckup
LLMCheckup is an interface for chatting with any LLM about its behavior. We connect several white-box and black-box interpretability tools (§3), s.t. LLMCheckup takes into account model internals, datasets and documentation for generating self-explanations. User requests for explanations are recognized via a text-to-SQL task performed by the LLM under investigation (§2.1-2.2).
We showcase a short dialogue between the user and LLMCheckup in Figure 1 and a longer dialogue featuring different operations in Appendix B. LLMCheckup can answer various questions related to the data as well as the model’s behavior. For example, in Figure 1 the user is interested in the rationale for a specific prediction and the model generates an explanation to justify the assigned label. LLMCheckup also suggests to have a look at another related operation (token-level importance scores) that can help explain model’s behavior (§2.4), but the user asks for a modified (augmented) version of the same instance instead. As a result, the model paraphrases the original question which can be then treated as a new sample and the user can further examine it by using the custom input functionality of LLMCheckup2.4).
@@ 2.1 System architecture
Figure 2 illustrates the interaction flow of LLMCheckup. When a user asks a question, it will be parsed as an SQL-like query by the LLM. E.g., the first user question in Figure 1 will be parsed as filter id 26 and rationalize. The corresponding parsed operation (i.e., filter and rationalize in our example, see Table 1 for the full list of operations) will then be matched and executed. For response generation, the explanation provided by the underlying interpretability method is converted into a natural language output using a template-based approach Slack et al. (2023); Feldhus et al. (2023) and is then displayed to the user.
Filter filter(id) Access single instance by its ID
includes(token) Filter instances by token occurrence
Prediction predict(instance) Get the prediction for the given instance
randompredict(number) Precompute a subset of instances at random
mistakes show|count(subset) Count or show incorrectly predicted instances
Data show(list) Showcase a list of instances
countdata(list) Count number of instances in the dataset
label(dataset) Describe the label distribution across the dataset
Meta data() Information related to the dataset
model() Metadata of the model
About function() Inform about the functionality of the system
self() Self-introduction of LLMCheckup
qatutorial(op_name) Provide explanation for the supported operations (tutorial)
Explain nlpattribute(inst., topk, method_name) Provide feature attribution scores
rationalize(instance) Explain the output in natural language
NLU keywords() Show common keywords in the data
similarity(instance, number) Output top $k$ similar instances in the dataset
Pert. cfe(instance) Generate counterfactuals
augment(instance) Augment the input text
Logic and(op1, op2) Concatenation of multiple operations
or(op1, op2) Selection of multiple filters
Table 1: All operations (mappings between a partial SQL-type query and a function) facilitated by LLMCheckup, including all explainability methods mentioned in §3 and other supplementary operations. Operations marked with () support the use of custom inputs (see more details in App. A).
2.2 Parsing
To recognize users’ intents, the deployed LLM transforms a user utterance into a SQL-like query. The SQL-based approach is needed to formally represent the available operations (see Table 1) and their “semantics” including all necessary attributes. For user intent recognition, we employ two methods: Guided Decoding and Multi-prompt Parsing.
2.2.1 Guided Decoding
@@
Guided Decoding (GD) ensures that the generated output adheres to predefined grammatical rules and constraints Shin et al. (2021) and that parses of the user requests align with predefined operation sets Slack et al. (2023). GD is generally more suitable for smaller LMs, since in-context learning may encounter instability attributed to the fluctuations in the order of provided demonstrations, and the formats of prompts Ma et al. (2023).
2.2.2 Multi-prompt Parsing
@@
As an alternative to GD, we propose and implement a novel Multi-prompt Parsing (MP) approach. While GD pre-selects prompts based on the embedding similarity with user input, the model does not see all the available operations at once and the pre-selection may not include the examples for the actual operation required. With MP, we test whether showing all possible operations in a simplified format (i.e., without any attributes such as instance ID or number of samples) and then additionally prompting the model to fill in more fine-grained attributes can improve performance.
As a first step, MP queries the model about the main operation (see list of operations in Table 1). Next, depending on the chosen operation, MP selects the operation-specific prompts with 2-7 demonstrations [^2] The number of demonstrations depends on the difficulty of operation, e.g., how many attributes it may have. (user query and parsed outputs examples) to generate the full parses that may include several attributes. E.g., for the user input "What are the feature attributions for ID 42 based on the integrated gradients?", we start by generating nlpattribute and then augment the parse with the second prompt and transform it into filter id 42 and nlpattribute integrated_gradient.
Since the output of the model is not constrained, unlike in GD, in the MP setting we need to check whether the model’s output matches any of the available operations and if there is no exact match we employ SBERT [^3] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 to find the best match based on the embedding similarity. We also implement checks to avoid possible hallucinations, e.g., if the model outputs an ID that is not present in the input we remove it from the parser output. §5.1 evaluates the performance of both parsing approaches.
Figure 3: LLMCheckup interface with welcome message, free-text rationale and sample generator buttons. Expert XAI level and OPRO strategy are selected. For example multi-turn dialogues, see Table 5 and Table 6.
2.3 Interface
LLMCheckup provides a user interface (Figure 3) including a chat window to enter questions and settings on the right panel, including XAI expertise level selection, custom inputs, prompt editor and export functionality for the chat history. It is implemented in Flask and can be run as a Docker container. LLMCheckup provides a chat window Slack et al. (2023), a dataset viewer Feldhus et al. (2023), a custom input history viewer and question suggestions for different operations. Together, these UI elements facilitate dataset exploration and provide sample questions for all available operations to inspire users to come up with their own questions.
On the right side of the window, there is a Prompt Editor with different options for prompt modification (§3.2). The icons associated with each strategy describe them in detail, including the corresponding prompts that can be appended after the default system prompt.
@@ 2.4 Key features
Out of the box, we include five auto-regressive LLMs representative of the current state-of-the-art in open-source NLP (as indicated in the left column of Table 2) available through Hugging Face Transformers Wolf et al. (2020). The diverse choice of models demonstrates that our framework is generalizable and supports various Transformer-type models. While Falcon-1B Penedo et al. (2023) and Pythia-2.8B Biderman et al. (2023) are available for users with limited hardware resources (RAM/GPU), it is generally not recommended to use them due to their small model size, which may negatively affect performance and user perception. Llama2-7B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) are both mid-sized with 7B parameters, while Stable Beluga 2 Mukherjee et al. (2023) is a fine-tuned version of Llama2-70B. To facilitate the deployment of large models in a local environment, LLMCheckup offers support for various forms of LLMs. This includes quantized models through GPTQ Frantar et al. (2023), loading models in 4-bits with the assistance of bitsandbytes Dettmers et al. (2022), and the implementation of a peer-to-peer solution using Petals Borzunov et al. (2023), enabling efficient deployment on a custom-level GPU.
@@
To help non-experts get background knowledge in XAI, we introduce a tutorial functionality. It is based on prompting with different roles corresponding to levels of expertise in XAI (Figure 3) and enables us to provide tailored meta-explanations of supported operations to individuals. For example, at the beginner level, we add a system prompt hinting at the expertise: “As a NLP beginner, could you explain what data augmentation is?” (Figure 4). In such a way, all users can receive meta-explanations according to their expertise.
Customized inputs & prompts
In comparison to TalkToModel Slack et al. (2023), which was limited to three datasets, LLMCheckup offers users the freedom to enter custom inputs (e.g. modified original samples or even completely new data points, see the Custom Input box on the right panel in Figure 3), going beyond just querying instances from specific provided datasets. In addition, inspired by PromptSource Bach et al. (2022), a Prompt Editor (see Prompt modification section on the right panel in Figure 3) supports inserting both pre-defined and fully customized prompts, allowing the users to control how downstream tasks and rationalization (§3.2) are performed. All custom inputs are saved and can be inspected and reused later via a dedicated custom input history viewer.
Suggestion of follow-up questions
To guide the user through the conversation, we implemented a suggestions mode. The user receives suggestions for related operations that LLMCheckup can perform based on the dialogue context, e.g., if the user asks about the top $k$ attributed tokens for a specific sample, they will receive a suggestion to have a look at the generated rationales since both operations belong to the “Explanation" category also displayed in the user interface. The suggestions are grouped into several categories as specified in Table 1 (see Appendix F for more detail).
2.5 Add-on features
@@
Since LLMs may sometimes generate incorrect responses Welleck et al. (2020), LLMCheckup allows users to access information by conducting search through external knowledge bases, promoted by the integration of Google Search [^4] https://github.com/Nv7-GitHub/googlesearch (Figure 5). In particular, it provides an external link that contains information relevant to the input sample(s). Users can cross-reference the retrieved information with the provided explanations, thereby achieving a more comprehensive understanding.
@@
Motivated by Malandri et al. (2023), LLMCheckup not only accepts text input from users but also provides support for other modalities such as images and audio. To facilitate this, we integrate packages and models tailored to each modality. For optical character recognition (OCR), we use EasyOCR [^5] https://github.com/JaidedAI/EasyOCR. For audio recognition, we employ a lightweight fairseq S2T [^6] https://huggingface.co/facebook/s2t-small-librispeech-asr model Wang et al. (2020) trained on Automatic Speech Recognition (ASR).
@@ 3 NLP explainability tools
While we introduce each explainability method individually, these methods can be interconnected through follow-up questions from users or suggestions provided by LLMCheckup to preserve context. Table 5 and Table 6 show examples of explanations for each supported explainability method by LLMCheckup.
@@ 3.1 White-box
Feature attribution methods quantify the contribution of each input token towards the final outcome. In LLMCheckup, we deploy various auto-regressive models (§2.4), for which Inseq Sarti et al. (2023) is used to determine attribution scores. We support representative methods from Inseq, including Input x Gradient Simonyan et al. (2014), Attention Bahdanau et al. (2015), LIME Ribeiro et al. (2016), and Integrated Gradients Sundararajan et al. (2017) [^7] Details on the Inseq integration are described in App. C..
@@ 3.2 Black-box
Augmentation involves synthesizing new instances by replacing text spans of the input while preserving the semantic meaning and predicted outcomes Ross et al. (2022). Data augmentation can be achieved by LLM prompting with or without providing a few demonstrations Dai et al. (2023). Alternatively, NLPAug [^8] https://github.com/makcedward/nlpaug can be used to substitute input words with synonyms from WordNet Miller (1995). Augmented texts can offer valuable insights into model behavior on perturbation tasks and prediction differences between them and their original texts.
@@
Unlike data augmentation, counterfactuals manifest as input edits causing the predicted outcome to be different Wu et al. (2021); Chen et al. (2023). Counterfactuals are generated by prompting LLMs with manually crafted demonstrations.
@@ 4 Use cases
In this paper, we demonstrate the workflow of LLMCheckup on two typical NLP tasks: Fact checking and commonsense question answering. Figure 1 and Appendix B show sample dialogues where user asks questions regarding rationalization, data augmentation and other operations based on the ECQA data Aggarwal et al. (2021) for commonsense question answering. The LLMCheckup repository includes all the necessary configuration files for different LMs and our use cases. They can be easily adopted to many other downstream tasks, data and Transformer-type models, demonstrated in a tutorial which will be available with the camera-ready version of our repository.
@@ 4.1 Fact checking
The importance of fact checking has grown significantly due to the rapid dissemination of both accurate information and misinformation within the modern media ecosystem Guo et al. (2022). COVID-Fact Saakyan et al. (2021) is a fact-checking dataset that encompasses various claims, supporting evidence for those claims, and contradictory claims that have been debunked by the presented evidence.
@@ 4.2 Commonsense question answering
Unlike question answering, commonsense question answering (CQA) involves the utilization of background knowledge that may not be explicitly provided in the given context Ostermann et al. (2018). The challenge lies in effectively integrating a system’s comprehension of commonsense knowledge and leveraging it to provide accurate responses to questions. ECQA Aggarwal et al. (2021) is a dataset designed for CQA. Each instance in ECQA consists of a question, multiple answer choices, and a range of explanations. Positive explanations aim to provide support for the correct choice, while negative ones serve to refute incorrect choices. Additionally, free-text explanations are included as general natural language justifications.